OpenStudy (anonymous):

Which is equivalent to i22?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

i^22=i^(4*5 + 2)=i^2=-1 if you have any confusion divide the power by 4 then whatever remainder you get that value you can easily find out!!!

OpenStudy (zarkon):

I really wish people would stop writing things like the original post. if you want \(i^{22}\) then try writing i^(22)